OpenStudy (watchmath):

Calculate \[\lim_{n\to\infty}(\sqrt[n]{n}-1)^n\]

OpenStudy (anonymous):

\[\lim_{n\to\infty}(\sqrt[n]{n}-1)^n\] \[\lim_{n\to\infty}(n ^{\frac{1}{n}}-1)^n\] Try a the smaller limit of n^(1/n) \[\lim_{n\to\infty}n ^{\frac{1}{n}} =\infty^{0}\] Anything to the 0th power is 1, so this piece of the limit is slowly approaching 1. Therefore, you are left with: \[\lim_{n\to\infty}(1-1)^n = \lim_{n\to\infty}0^n = 0\] The answer is 0. I don't think I did a very good job of using limits though. http://www.wolframalpha.com/input/?i=Plot%5B%28n^%281%2Fn%29+-+1%29^n%2C+ {n%2C0%2C100}%5D